15-th Conference of the International Linear Algebra Society

February 19, 2008

Professor Tingzhu Huang School of Applied Mathematics University of Electronic Science and Technology of China Chengdu, Sichuan, 610054 P. R. China

Dear Professor Huang:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society invites you to participate in our Conference delivering a lecture about your recent research results.

The 15-th Conference will be held in Cancún, México, from June 15 to June 21, 2008.

Some useful information is available at the conference web-page http://star.izt.uam.mx/ILAS08

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially

Dr. Luis Verde-Star Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa Apartado 55-534, Mexico DF 09340, Mexico Tel. +52 55 5804 4961, Fax +52 55 5804 4660, e-mail: verde@star.izt.uam.mx